### v2

# SlumpGuard: An AI-Powered Real-Time System for Automated Concrete Slump Prediction via Video Analysis

Abstract
Concrete workability is essential for construction quality, with the slump test being the most common on-site method for its assessment. However, traditional slump testing is manual, time-consuming, and prone to inconsistency, limiting its applicability for real-time monitoring. To address these challenges, we propose SlumpGuard, an AI-powered, video-based system that automatically analyzes concrete flow from the truck chute to assess workability in real time. Our system enables full-batch inspection without manual intervention, improving both the accuracy and efficiency of quality control. We present the system design, the construction of a dedicated dataset, and empirical results from real-world deployment, demonstrating the effectiveness of SlumpGuard as a practical solution for modern concrete quality assurance.
keywords:
Automated quality control , Chute detection , Optical flow , Slump prediction , Video classification
journal: Automation in Construction
[…]
[label1]organization=Yonsei University,addressline=50 Yonsei-ro, Seodaemun-gu, city=Seoul, postcode=03772, country=Republic of Korea
[…]
[label2]organization=GS Engineering & Construction Corp, addressline=33 Jong-ro, Jongno-gu, city=Seoul, postcode=03159, country=Republic of Korea
[…]
[label3]organization=Seoul National University, addressline=1, Gwanak-ro, Gwanak-gu, city=Seoul, postcode=08826, country=Republic of Korea
1 Introduction
[…]
Concrete stands as the backbone of modern construction infrastructure, with over 14 billion cubic meters produced annually worldwide, making it the most widely consumed material after water [1]. Among the critical properties governing concrete performance, workability—primarily assessed through slump testing—represents a fundamental parameter that directly influences mixing efficiency, transportation logistics, placement operations, and compaction effectiveness on construction sites [2, 3]. However, the complex and dynamic nature of construction environments presents unprecedented challenges for maintaining consistent concrete quality control. In particular, conventional manual slump testing, still the industry standard, often fails to meet the demanding requirements of large-scale, time-sensitive projects.
The construction industry therefore continues to struggle with real-time quality assurance. Traditional slump testing remains labor-intensive, operator-dependent, and unable to provide continuous monitoring during concrete placement. Manual testing introduces delays, increases costs, and is sensitive to environmental factors, while reliance on less experienced personnel further undermines consistency [4]. Although automated approaches such as video-based slump prediction, stereo vision, depth sensing, and IoT monitoring have been explored, most are validated only in controlled or small-scale settings and face challenges such as variable lighting, dust, and equipment interference in dynamic construction environments [5, 3, 6]. Moreover, these systems often depend on specialized hardware or require significant infrastructure modifications, limiting their practicality. Consequently, there is an urgent industry need for field-deployable, cost-effective solutions capable of providing robust, real-time monitoring of concrete workability without disrupting construction workflows, addressing environmental variability, and integrating seamlessly with existing operations [7, 8].
To address the limitations of conventional slump testing in real-site applications, we present SlumpGuard, a comprehensive AI-powered system specifically engineered for real-world deployment in construction environments. Unlike manual methods that are labor-intensive, inconsistent, and unsuitable for continuous monitoring, SlumpGuard operates directly on the flow of ready-mix concrete from mixer truck chutes to enable automated, full-batch inspection of every delivery without manual intervention. Our approach introduces several key innovations: (1) a robust three-stage pipeline combining oriented object detection, optical flow analysis, and video classification, (2) advanced data augmentation strategies specifically designed to handle the environmental variability encountered in construction sites, and (3) comprehensive validation using a dataset collected over eight months from actual construction conditions rather than laboratory simulations. This video-based system not only enhances the consistency and reliability of slump measurements but also streamlines quality assurance procedures across the entire construction workflow, ultimately contributing to safer, more efficient, and higher-quality construction outcomes.
The remainder of this paper is organized as follows. Section 2 reviews existing studies on concrete workability assessment and automated monitoring approaches. Section 3 describes the dataset collection process and annotation methodology. Section 4 details the architecture of the proposed SlumpGuard system, including its three-stage pipeline and data augmentation strategies. Section 5 presents experimental results and performance evaluation under real-site conditions. Finally, Section 6 concludes the paper and outlines future research directions.
2 Background and Related Studies
2.1 Optical Flow
Optical flow is a computer vision technique used to estimate the motion of objects or scenes across video frames by analyzing the movement of pixels. It represents motion as a vector field, where each vector indicates the displacement of a pixel between consecutive frames. There are two main approaches: dense optical flow, which computes motion for every pixel, and sparse optical flow, which estimates motion only at selected feature points. While dense optical flow [9] offers a detailed view of the entire scene’s motion, it is computationally intensive. Sparse optical flow is more efficient but may overlook fine-grained motion.
In this work, we adopt the Lucas-Kanade method [10], a widely used algorithm for sparse optical flow, suitable for real-time applications such as tracking concrete flow. It assumes that the intensity of a point remains constant over time:
$I(x,y,t)=I(x+\Delta x,y+\Delta y,t+\Delta t),$ (1)
and linearizes this relation using a Taylor series expansion, yielding:
$\frac{\partial I}{\partial x}u+\frac{\partial I}{\partial y}v+\frac{\partial I}{\partial t}=0,$ (2)
where $u$ and $v$ are the horizontal and vertical components of the flow. By solving this equation over a small window around each feature point, the Lucas-Kanade method efficiently estimates the local motion.
However, classical optical flow methods, including Lucas-Kanade, often struggle under challenging conditions such as illumination changes, occlusions, or non-rigid motion. To overcome these limitations, deep learning-based methods have emerged as the new state-of-the-art.
FlowNet [11] pioneered the use of convolutional neural networks for end-to-end optical flow prediction, learning motion patterns directly from synthetic datasets. Building on this idea, FlowNet2 [12], PWC-Net [13], and LiteFlowNet [14] introduced architectural improvements such as pyramid-based warping, cost volume reasoning, and hierarchical refinement. More recently, RAFT [15] achieved significant performance gains by leveraging dense, all-pairs correlations and iterative updates to produce highly accurate flow estimates.
Previous studies have applied optical flow combined with imaging techniques for slump prediction under controlled experimental settings [4, 5]. In contrast, our approach is specifically designed for robust, real-world deployment on construction sites, enabling continuous, automated slump quality monitoring under diverse and dynamic field conditions.
2.2 Object Detection
Object detection is a fundamental computer vision task widely applied in construction to automatically identify and localize key entities such as workers, equipment, and materials in images or videos. Recent studies have demonstrated its effectiveness in construction site monitoring [16, 17].
State-of-the-art detection frameworks such as Faster R-CNN [18], Mask R-CNN [19], SSD [20], and YOLO [21] offer a good balance between speed and accuracy, making them well-suited for real-time applications on construction sites.
However, in construction environments where objects like materials and equipment often appear at arbitrary angles, standard axis-aligned bounding boxes may be insufficient. In such cases, oriented object detection becomes essential. Early approaches like R2CNN [22] introduced rotation-invariant proposals, while more recent models such as Oriented R-CNN [23] and ReDet [24] employ rotation-aware architectures to improve detection accuracy. These methods are particularly effective for automated inspection tasks involving rotated or non-axis-aligned objects.
2.3 Video Classification
Video classification is a fundamental task in computer vision that involves assigning semantic labels to video clips by analyzing both spatial and temporal information inherent in the data. Recent advances in this domain have been largely driven by two primary classes of deep learning architectures: 3D convolutional neural networks (3D CNNs) and vision transformers [25] tailored for video.
3D CNNs like ResNet3D [26] extend 2D convolutions to the temporal dimension, effectively capturing spatiotemporal features and achieving strong results on benchmarks such as Kinetics-400 [27]. Transformer-based models, notably TimeSformer [28], use divided spatial and temporal self-attention to model long-range dependencies efficiently, surpassing many convolutional approaches. Further advancements include VideoSwin [29] and MViT [30], which incorporate hierarchical and multiscale attention mechanisms to improve video representation learning and classification performance.
These video classification models have also been increasingly applied in the construction industry, playing a pivotal role in streamlining quality supervision, activity recognition, and concrete workability assessment. Previous studies have also explored video-based approaches for material quality monitoring and evaluation [31, 3, 8], alongside behavior classification systems aimed at improving worker safety and productivity [32, 33].
3 Dataset
3.1 Data Collection
Figure 1: An example of our collected dataset. The concrete pouring from two concrete mixer truck’s chutes was captured using a ZED 2i stereo camera. The overall resolution of each camera system is $3840\times 1080$, consisting of two individual cameras with $1920\times 1080$ resolution each.
In this study, we constructed a high-quality video dataset to accurately capture the dynamic slump behaviors of ready-mixed concrete during concrete placing. To achieve this, we conducted controlled concrete pouring experiments using two ready-mix concrete trucks, each carrying concrete mixtures with predetermined slump values.
In the experimental process, considering cost efficiency, we strategically utilized a single stereo camera for video acquisition. The stereo camera allowed simultaneous capturing of video footage from two distinct viewpoints in a single filming session, allowing richer data collection within a limited number of experiments. This approach maximized data efficiency while clearly capturing the concrete flow through the chute. Additionally, we performed traditional slump cone measurements at regular intervals, establishing precise ground-truth labels corresponding to each video frame. This example is shown in Figure 1.
3.2 Preprocessing and Annotation
Since we employed a stereo camera setup, the collected videos consisted of two separate streams. Therefore, as part of preprocessing, the original stereo data was divided into two individual video sequences, effectively doubling the amount of data for analysis. Considering that concrete discharges directly from the ready-mixed concrete chute, our preprocessing specifically targeted the chute region. Bounding box annotations were performed by categorizing into two classes, as shown in Figure 2, an “Unrotated Chute” representing the entire chute region with axis-aligned boxes, and a “Chute” represented by rotated bounding boxes.
Figure 2: An example of data annotation for detecting concrete pouring regions using bounding boxes. The red boxes represent the chutes, while the blue boxes indicate the corresponding unrotated bounding boxes.
4 Automation Strategy
Figure 3: Overview of our pipeline. Our pipeline consists of three steps: detecting chutes, identifying when and from which chute an object falls, and predicting the slump. Finally, we compare the predicted slump with the requested range to perform anomaly detection. In the figure, “URChute” refers to an unrotated chute. The variables $x,y,w,h$, and $\theta$ represent the bounding box’s center $x$-coordinates, $y$-coordinates, width, and height, respectively, with $\theta$ measured in degrees.
To practically implement our pipeline at construction sites, we propose automating the entire pipeline. Specifically, the following three aspects will be automated:
  1. 1.
    Determining from which of the two chutes the concrete is discharged.
  2. 2.
    Identify the exact time when the concrete begins to flow from the chute.
  3. 3.
    Assessing the discharged concrete slump to determine whether it falls within a specified range.
An overview of our proposed automation pipeline is presented in Figure 3.
4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model. We employ YOLOv8 [34], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [35], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated-bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
To enhance system efficiency, if the object detector continuously identifies the same region for more than $8$ frames, we halt the model’s inference and fix the Region of Interest (RoI) in the video. Specifically, the RoI is determined by averaging the bounding box coordinates over the detected frames, ensuring stability in localization. This approach reduces GPU usage for every frame, thereby improving power efficiency while maintaining detection accuracy.
To focus on the chute interior, we convert the detected rotated bounding box into an upright rectangular patch before cropping. The resulting coordinates are then used to crop the unroated chute area, generating separate videos for the left section, $V^{L}$, and the right section, $V^{R}$, of the chute. By using these segmented videos, we extract only the information related to the concrete inside the chute, allowing the model to better understand the concrete features.
4.2 Detecting Concrete Placement Location and Timing
On construction sites, concrete is discharged from two separate chutes connected to individual concrete mixer trucks. In this study, we detect which chute the concrete is falling from and accurately identify the starting point of the flow. This step is essential to ensure that the subsequent video classification model receives only the segments where concrete is actually flowing, enabling more precise slump analysis. Since video segments without concrete flow are not relevant for analysis, filtering out such irrelevant data in advance is a critical preprocessing step.
To precisely determine when and from which chute the concrete starts to fall, we perform motion analysis based on the segmented chute videos $V^{L}$ and $V^{R}$ obtained in Step 1. For each video segment, we extract bounding boxes for the chutes from the object detector and compute optical flow at the center point of each bounding box. Specifically, we apply the Lucas-Kanade method, a sparse optical flow algorithm, to track the motion vector $(u,v)$ at the center $(x,y)$ across consecutive frames. When the tracked center point crosses the bottom boundary of the bounding box over time, we identify the moment of concrete drop as well as the specific chute through which the concrete is discharged.
To determine whether the center point $(x_{t},y_{t})$ passes through the bottom edge of a rotated bounding box at the time $t$, we proceed as follows:
4.2.1 Definition of the Bottom Edge Direction
Given a rotated bounding box coordinates, the bottom edge lies perpendicular to the vertical axis of the box. The direction vector of the bottom edge is defined as:
$\mathbf{d}_{bottom}=\begin{bmatrix}\cos(\hat{\theta})\\ \sin(\hat{\theta})\end{bmatrix}\quad\text{where, }\hat{\theta}=\frac{\theta\pi}{180}.$ (3)
Note that $\theta$ is converted from degrees to radians, $\hat{\theta}$, since trigonometric functions are defined in radians. Next, we compute two endpoints $\mathbf{p}_{1}$ and $\mathbf{p_{2}}$ of the bottom edge by shifting from the center $(x,y)$ along the direction of the horizontal axis (which is orthogonal to the vertical axis) by half of the width:
[…]
4.2.2 Determining the Location and Timing
To make the temporal change explicit, we define the center point of the tracked chute in each frame as $(x_{t},y_{t})$ at time $t$, and as $(x_{t-1},y_{t-1})$ at the previous frame t1t-1. The optical flow vector $(u,v)$ represents the displacement from frame t1t-1 to $t$, so that $(x_{t},y_{t})=(x_{t-1}+u,y_{t-1}+v)$. The endpoints of the bottom edge of the rotated bounding box, $\mathbf{p_{1}}=(x_{1},y_{1})$ and $\mathbf{p_{2}}=(x_{2},y_{2})$, are computed.
To determine whether the tracked center point crosses the bottom edge between frames t1t-1 and $t$, we explicitly check the positions at both time steps relative to the edge. First, the line equation for the bottom edge is:
$m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}+\epsilon},\quad b=y_{1}-mx_{1},$ (5)
where $\epsilon$ is a small constant to avoid division by zero. The signed vertical distance from the center point to the bottom edge at each time step is given by:
$d_{t}=y_{t}-(mx_{t}+b),\quad d_{t-1}=y_{t-1}-(mx_{t-1}+b).$ (6)
The center point is considered to have crossed the bottom edge if the signs of $d_{t}$ and $d_{t-1}$ are difference or zero:
$d_{t}\times d_{t-1}\leq 0.$ (7)
This approach allows us to identify both the active discharge chute and the timing of the concrete placement.
4.3 Slump Prediction and Anomaly Detection
To enable efficient and real-time estimation of concrete slump, we leverage the drop timing and location identified in Section 4.2. Specifically, given the drop time $t$, we predict the slump using video data from the corresponding chute region. To focus on the continuous flow of concrete, we adopt a video classification model, ResNet 3D [26]. Let $V$ denote the chute video (i.e., $V^{R}$ or $V^{L}$), and $t$ the detected drop time.
4.3.1 Data Preprocessing
We crop the unrotated-bounding box corresponding to $V$ from the full image to isolate the chute area. From the frame at time $t$, we extract a sequence of $N$ consecutive frames from the cropped region to construct a temporally continuous input.
4.3.2 Data Augmentation
Given that construction sites are exposed to a wide range of lighting and environmental conditions, data augmentation is essential to improve the model’s generalization performance. To simulate these real-world variations, we employ a set of augmentation techniques, including ColorJitter, gamma correction, horizontal flipping, and contrast adjustment. We also apply MixUp [36] and CutMix [37], which blend either the pixel content or the labels of two training examples to generate mixed samples. While MixUp interpolates both images and labels, CutMix replaces a region of an image with a patch from another image, adjusting the labels proportionally to the area. These strategies encourage the model to learn more robust and smoother decision boundaries, and to be less sensitive to spurious correlations between local image regions and class labels.
Together, these augmentations help the model handle diverse visual appearances–such as changes in brightness, shadows, occlusion, and orientation–that are commonly encountered in construction sites environment. They also promote better regularization by expanding the training distribution and reducing overfitting, ultimately leading to improved performance under unseen conditions.
4.3.3 Model
We employed the ResNet 3D architecture [26] as our video classification backbone, designed to jointly capture spatial and temporal patterns. We experimented with two variants: Mixed Convolution and (2+1)D convolution. The Mixed Convolution model uses 3D convolutions in early layers for motion modeling and 2D convolutions in later layers for efficient spatial reasoning, achieving performance comparable to full 3D models with fewer parameters. The (2+1)D model factorizes 3D convolutions into separate spatial and temporal steps, enhancing optimization and model capacity.
4.3.4 Training
To formulate the slump estimation task as a classification problem, we categorized the continuous slump values into five categorical intervals (e.g. 150–180mm, 180–210mm), where the width of each interval reflects the $\pm 30$ mm tolerance specified in the KS F 4009 standard [38], along with an additional margin for human error. For intervals at the extremes (i.e. below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
To enhance generalization, we applied the MixUp technique, which linearly interpolates both input videos and their corresponding labels to generate augmented samples with soft targets. Under this setting, we replaced the standard cross-entropy loss with a soft-target formulation that incorporates label smoothing:
$L_{cls}=-\lambda_{cls}\sum_{i=1}^{C}y_{i}^{soft}\log(\hat{y}_{i}),$ (8)
where $y_{i}^{soft}\in[0,1]$ denotes the soft label vector obtained via MixUp and label smoothing, and $\hat{y}_{i}$ represents the predicted probability for class $i$.
By leveraging softened target distributions and encouraging smoother decision boundaries, this loss formulation helps mitigate overfitting to minority classes and reduces bias toward dominant categories—issues that frequently arise in imbalanced datasets. Consequently, the model is guided to generalize more effectively across all classes, regardless of their frequency in the training distribution.
4.3.5 Slump Prediction and Anomaly Detection
We use the trained video classification model to predict the slump of concrete. To enhance both real-time prediction and robustness, we perform the prediction $T$ times using $N$ consecutive frames each time, and determine the final result via majority voting. This allows for a more stable and reliable estimation of the slump. Based on the predicted slump category, we compare it with the ordered slump range at the construction site. If the prediction falls within the specified range, the concrete is considered acceptable; otherwise, it is flagged as abnormal, enabling effective quality control of the material.
5 Evaluation of Prediction Models
We perform both quantitative and qualitative evaluations at each step of the system to thoroughly assess the effectiveness of each component. Specifically, we analyze how each module contributes to overall performance through ablation studies and visualization-based inspections.
Figure 4: Statistics of our dataset. The outer pie shows the distribution of slump ranges, while the inner pie details the specific slump values within each category.
5.1 Dataset
Over an eight-month period, a dataset was constructed by replicating real-world construction site conditions in a controlled experimental settings. The recorded videos were segmented into 10-second intervals, resulting in $4,504$ training samples, $998$ validation samples, and $941$ test samples, as shown in Figure 4.
5.2 Chute Detection
5.2.1 Evaluation Metrics
To evaluate the performance of chute detection, we adopt evaluation metrics commonly used in object detection tasks, namely mAP50-95 and precision. Specifically, the mean average precision (mAP50-95) is computed as the average of AP values at IoU thresholds ranging from 0.50 to 0.95 in increments of 0.05:
$\text{mAP}_{50-95}=\frac{1}{10}\sum^{0.95}_{\text{IoU}=0.50}\text{AP}_{\text{IoU}}$ (9)
This metric provides a comprehensive measure of the detection accuracy at varying levels of localization strictness. Additionally, we report Precision, which quantifies the proportion of correctly predicted positive instances among all predicted positives:
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (10)
where TP and FP denote true positives and false positives, respectively. Together, these metrics offer a robust assessment of both the detection quality and reliability of the chute localization module.
5.2.2 Results
Our chute detection model achieved strong quantitative results, with an mAP50-95 of 0.9945 and a precision of 0.995. The qualitative results are shown in Figure 5, where all bounding boxes exhibit confidence scores above $0.95$, indicating highly reliable detections. This demonstrates the strong and effective performance of our chute detection model. Furthermore, real-time inference is feasible, as the model achieves an average speed of 9.09.0ms per frame, making it suitable for real-time applications.
Figure 5: Qualitative results of chute detection. Blue and Sky blue bounding boxes represent detections of “URChute” and “Chute”, respectively. The number inside each box denotes the confidence score.
5.3 Detecting Concrete Placement Location
Under 150 150 $\sim$180 180 $\sim$210 210 $\sim$240 Over 240 Avg
Left 100 100 100 100 98.45 99.91
Right 98.60 95.00 91.90 88.70 95.19 92.89
None 100 100 100 100 100 100
Avg 99.30 97.68 95.40 94.51 96.92
Table 1: Results of concrete placement location across various slump conditions and chute position. “None” means no pouring situation. All values in the table are in accuracy (%).
To evaluate which chute of the mixer truck the concrete was poured from, human annotators manually labeled the ground truth for each instance. Identifying the chute from which the concrete is poured also allows for accurate determination of the pouring start time, thereby enabling more precise quantitative evaluation. To facilitate this evaluation, we conducted experiments under three scenarios: pouring from the left-side chute, pouring from the right-side chute, and no pouring. This result is shown in Table 1.
As shown in the Table 1, the proposed system demonstrated an overall accuracy exceeding $90\%$ on average, and achieved $100\%$ accuracy in cases where concrete placement did not occur. While the chute located on the left side exhibited $100\%$ accuracy in most experimental cases, several errors were observed for the chute on the right side. This was primarily attributed to the shadow cast by the hopper cover located above the right-side chute. In particular, it was observed that after approximately 10 seconds from the start of the video, the direction of the optical flow vectors $(u,v)$ tended to follow the movement of the shadow rather than the actual flow of the concrete. However, such errors are not expected to pose significant issues in practical field applications when determining whether concrete is being placed. Additionally, in samples with slump values exceeding 240 mm, a temporary decline in accuracy was also observed for the left-side chute. This phenomenon occurred when the slump was excessively high, resulting in smooth and fluid concrete flow, which made pixel-level motion detection using optical flow more difficult. This issue was observed only in the sample with a slump value of 280 mm.
These experimental results demonstrate that the proposed system can accurately detect active chutes with high reliability. Since chute identification directly corresponds to detecting the start of concrete placement, the system is proven to be effective for automatic recognition of the placement zone and estimation of the starting time in real-world construction environments.
5.4 Slump prediction
5.4.1 Model Training Setup
We trained the model using 16-frame video clips, sampled at a frame interval of 2, and extended by a temporal factor of 2. The resize ratio was randomly selected from { 3/4, 4/3 }, and standard data augmentations were applied. For spatial augmentation, we used random horizontal flipping with a probability of 0.5 and ColorJitter with the following ranges: brightness 0.4, contrast 0.4, saturation 0.4, and hue 0.1. Additionally, MixUp ($\alpha$ = 0.2, probability = 0.5) and CutMix ($\alpha$ = 0.2, probability = 0.5) were employed to generate soft-label training samples, encouraging better generalization and robustness to noisy data. Label smoothing with a factor, $\lambda_{cls}$, of 0.1 was also applied.
The model was optimized using the AdamW optimizer [39] with a learning rate of 1e41e-4, and a weight decay of 1e41e-4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [40] with cosine annealing, with a peak learning rate set to 1e31e-3.
Val Test
Backbone Architecture Acc F1 Acc F1
Transformer TimeSFormer [28] 0.6980 0.7280 0.5724 0.6155
ResNet 3D MC3-18 0.8080 0.8069 0.7916 0.8526
R3D-18 0.8152 0.8183 0.8011 0.8511
R(2+1)D-18 0.8152 0.8149 0.8226 0.8691
Table 2: Evaluation results across various video classification model architectures.
Label smoothing Weighted sample MixUp Acc F1
0.7654 0.8180
0.7717 0.8205
0.7955 0.8499
0.8226 0.8691
Table 3: Ablation study results on the training strategy for R(2+1)D-18. “✓” indicates that the technique was applied, while a blank cell denotes that it was not.
5.4.2 Results
We evaluate our model’s ability to predict the range of concrete slump using both the validation and test sets. Accuracy and F1-score are used as the primary evaluation metrics. In addition to evaluating the overall performance, we conduct an ablation study to assess the effectiveness and robustness of each component of our proposed model.
As shown in Table 3, our models employing ResNet-3D as the backbone achieve over 80% accuracy on both the validation and test sets. In contrast, the Transformer [41]-based TimeSFormer [28] exhibits lower accuracy. Among the evaluated architectures, we selected R(2+1)D-18 as our final model, as it achieved the highest performance on both validation and test sets. Based on this architecture, we conducted a detailed ablation study, and the corresponding results are reported in Table 3.
To mitigate data imbalance and prevent overfitting, we incorporated label smoothing, weighted sampling, and MixUp-based data augmentation. Each of these techniques led to measurable performance gains, as demonstrated in our experiments. Through these experiments, we successfully developed an optimized and robust model for concrete slump prediction.
6 Conclusion
This paper presented SlumpGuard, an automated system for monitoring and classifying concrete slump using video analysis. The proposed approach employs a three-stage pipeline, consisting of chute detection, pouring location and timing estimation, and concrete slump classification. Specifically, a YOLO-based object detection model was developed to identify the concrete mixer truck and its discharge chute. Based on these detections, optical flow analysis was employed to accurately determine the pouring location and timing. Finally, a video classification model was used to predict the concrete slump, enabling automated and comprehensive inspection of all incoming concrete deliveries, thereby replacing manual sampling procedures and contributing to more reliable quality assurance in concrete construction. Conclusions are drawn as follows.
  1. 1.
    To facilitate video-based concrete slump prediction, we constructed a dataset comprising 6,443 video clips, each 10 seconds long, capturing various concrete pouring scenes. Bounding box annotations for the discharge chute of the mixer truck were provided to enable the development of deep learning-based methods.
  2. 2.
    The proposed YOLO-based chute detector demonstrated highly reliable performance in identifying the discharge chute of the mixer truck. It achieved a mean Average Precision (mAP50:95) of 99.45 and a Precision of 99.5, indicating a high level of accuracy.
  3. 3.
    We developed an algorithm to determine the location and timing of concrete discharge using the computer vision technique of optical flow. Experimental results demonstrated an average accuracy of over 95%.
  4. 4.
    We showed that the concrete slump range can be automatically predicted using a video classification model. Through a series of comparative experiments and ablation studies, we identified the optimal model architecture and hyper-parameters, achieving a high-accuracy slump prediction model with a performance of 82%. Based on these results, our system can automatically verify whether the predicted slump range matches the ordered specification, enabling the development of an automated quality control system for concrete.
Future work could focus on predicting continuous concrete slump values rather than categorical ranges. In addition, based on the results of this study, the system could be integrated with Building Information Modeling (BIM) or construction process management systems to enable automated tracking of concrete quality throughout the entire construction lifecycle, ultimately leading to a more practical and deployable solution.
Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea Government (MSIT) (No. RS-2024-00354218, No. RS-2020-II201361, Artificial Intelligence Graduate School Program at Yonsei University), and by the Ministry of Land, Infrastructure and Transport of the Korea Government through the Railway Technology Research Program (No. RS-2021-KA163289).
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.